# Misaligning Reasoning with Answers - A Framework for Assessing LLM CoT Robustness

### v1

@@ 5.1 Experimental setup
Datasets. We use three datasets in line with existing literature on reasoning (wei2022chain, ; trivedi2022interleaving, ; kojima2022zerocot, ; diao2023active, ). For math reasoning, we select two math datasets SingleEq (koncel2016mawps, ) with single-step math problems and GSM8K (cobbe2021training, ) with multi-step math problems. Apart from that, we select StrategyQA (geva2021did, ) for commonsense reasoning. For all datasets, we use the test split for the evaluation. Specifically, we sampled $500$ questions for the evaluation of the StrategyQA dataset and used all data (508 for SingleEq and 1319 for GSM8K) in the other two datasets. More details about the datasets can be found in Appendix A.2.
Models. Our experiments focus on open-source and closed-source models for robustness evaluation. For open-source models, we use Llama-3-8B (dubey2024llama, ), Mistral-7B (jiang2023mistral, ), Zephyr-7B-beta (tunstall2023zephyr, ), Qwen2.5-7B (yang2024qwen2, ), and DeepSeek-R1-Distill-Qwen-7B (guo2025deepseek, ). We evaluate the transferability of white-box pertrubations to closed-source models GPT-4 and GPT-3.5-turbo (gpt4, ). Also, we use GPT-3.5-turbo as our judging model for comparing COTs, and DeepSeek-R1-Distill-Qwen-7B (guo2025deepseek, ) for evaluating the semantic meanings of perturbed questions.
@@
Implementation. For token-level perturbation on open-source models, we set the inserted ratio to be $a=0.2$ for all datasets; for the replacement ratio $k$, we set it to be $(0.5,0.25,0.5)$ for SingleEq, GSM8K, and StrategyQA datasets, respectively. For embedding-level perturbation, we choose hyperparameters with perturbation percentage $\epsilon=(0.02,0.005,0.03)$ for SingleEq, GSM8K, and StrategyQA datasets, which are imperceptible for the token mapping. Also, we set the number of perturbation steps $j$ to be 5 for all datasets and two kinds of perturbations. For closed-source model experiments, we set temperature $T=0.7$ and we test the closed-source models on the adversarial examples that open-source models generate. The default version of GPT-3.5-turbo used is GPT-3.5-turbo-0125, and the GPT-4o used is GPT-4o-2024-08-06. Apart from that, we set the maximum length of generating output to be 256. We use a conventional ICL setting for our experiments (brown2020language, ; rae2021scaling, ). We apply the same number of exemplars as wei2022chain and use the exemplars from diao2023active , which are 8 for SingleEq and GSM8K, as well as 6 for StrategyQA. Runtime, LLM judge analysis, and limitations can be found in Appendix A.5 and A.7.
@@ 5.2 Comparison with baselines
In Table 3, we compare MATCHA with the following baselines for token-level perturbations: (1) Random: we randomly insert tokens into the original question; (2) GCG: we optimize a suffix similar to gcg using our defined loss; (3,4,5) MATCHA: we experiment with MATCHA with random locations/with only CoT loss/full versions. We show that MATCHA can achieve a comparatively lower unattackable rate and higher success rate, indicating MATCHA token-level perturbations have stronger perturbation effectiveness.
@@ 5.4 Transferability to closed-source models
Dataset-wise transferability: Transferability rates are usually lower for SingleEq compared to GSM8K, suggesting that simpler math problems with single-step reasoning are harder to attack. Since these problems involve fewer reasoning steps, there are fewer opportunities for perturbations to disrupt the model’s logical progression. In contrast, multi-step reasoning tasks (GSM8K) show higher transfer rates, indicating that perturbing intermediate steps in complex problem-solving is an effective attack strategy. Also, for commonsense reasoning, the results show a moderate transferability.
Comparison between GPT-3.5-turbo and GPT-4o: Overall, GPT-4o exhibits comparatively similar transferability rates compared to GPT-3.5-turbo, suggesting that input perturbations may not yet be sufficiently investigated or incorporated during training, even in more advanced models.
@@ 5.5 Examples of successful attacks
Wrong calculations refers to the errors in the calculation steps. As shown in the first row of Figure 3, the model thinks "20/3=5", which is incorrect in the calculation. Wrong reasoning refers to the evident reasoning errors in the procedure of arriving at the correct solution. For the second example in Figure 3, the model makes the mistake of getting the total number of running hours (should be $3+2/3*2=6$ instead of $4.5$), leading to the overall error in the reasoning procedures. Wrong information refers to the model outputting some information that is wrong according to the question, to arrive at the correct answer. We present an example in the third column of Figure 3, where the model makes up "Sophocles" instead of "Sophist" to arrive at the correct answer. Unrelated information refers to the model outputting some unrelated information when trying to answer the question. In the fourth column of Figure 3, the model starts to make up non-existent information like "There are 15 trees in the grove" to guess the answer. The complete list of successful attack examples can be found in Appendix A.6.
@@ 6 Conclusion
We introduce MATCHA, a novel evaluation framework, to show the misalignment between reasoning/answer and evaluate the robustness of LLM reasoning. Our results show that LLMs are highly sensitive to input perturbations, especially in multi-step and commonsense reasoning. Our black-box transferability experiments also reveal a notable attack success rate on closed-source models like GPT, highlighting broader vulnerabilities. These findings underscore the need for consistency between answer and reasoning, as well as more training strategies to enhance the reliability of CoTs.
@@ Appendix A Appendix
1:Number of attack steps $k$, ICL examples $I$ with length $n_{1}$, unlabeled question $x_{p}$ with length $n_{2}$, ground truth $y_{gt}$, model $f$ parameterized by $\theta$, attack region $\epsilon$, stepsize $\alpha$, loss weight $\lambda$, embedding layer $h$, judge LLM $J$.
@@
4$\triangleright$ Initialization for conditions
5:$(y_{p},r_{p})\leftarrow\operatorname*{argmax}P_{\theta}((y_{p},r_{p})|I,(x_{p}))$
6$\triangleright$ Generate original answer and CoT
@@
10:     $E_{a}\leftarrow h(I,(x_{p},y_{p}))$triangleright Get embeddings
11:     $q_{a},q_{c}\leftarrow f(E_{a}),f(E_{c})$triangleright Get logits
@@
16:         $\mathcal{L}_{opt}\leftarrow\mathcal{L}_{c}-\lambda\cdot\mathcal{L}_{a}$triangleright Loss objective
@@
19:         $E^{\prime}_{c}\leftarrow\text{concat}(E_{c}[:n_{1}],E_{tmp}[n_{1}:n_{2}],E_{c}[n_{2}:])$triangleright Update perturbed embedding
@@
33:return $a,b,c,d$triangleright Final conditions
@@
1:Number of attack steps $k$, ICL examples $I$ with length $n_{1}$, input ids and input tokens, unlabeled question $x_{p}$ with length $n_{2}$, ground truth $y_{gt}$, model $f$ parameterized by $\theta$, embedding layer $h$, selected number of token $T_{k}$, judge LLM $J_{1}$ for reasoning correctness and $J_{2}$ for question correctness.
@@
4$\triangleright$ Initialization for conditions
5:$(y_{p},r_{p})\leftarrow\operatorname*{argmax}P_{\theta}((y_{p},r_{p})|I,(x_{p}))$
6$\triangleright$ Generate original answer and CoT
@@
9:     $q_{a},q_{c}\leftarrow f(I,(x_{p},y_{p})),f(I,(x_{p},y_{p},r_{p}))$triangleright Get target logits
@@
11:     $x^{\prime}_{r}\leftarrow\text{random\_insert}(x^{\prime}_{p})$triangleright Randomly insert tokens into the question part
@@
[…]
@@
22:         $q^{\prime}_{c},q^{\prime}_{a}\leftarrow f(I,(x_{p},y_{p})),f(I,(x^{\prime}_{r},y_{p},r^{\prime}_{p}))$triangleright Get perturbed logits
@@
24:         $\mathcal{L}_{opt}\leftarrow\mathcal{L}_{c}-\lambda\cdot\mathcal{L}_{a}$triangleright Loss objective
25:         $g_{tok}\leftarrow\nabla_{e}(-\mathcal{L}_{opt})$triangleright $g_{tok}$ has a shape of [seq_len, vocal_size]
[…]
28:         $s_{tok}\leftarrow\text{Topk}(s_{tok},T_{k})$triangleright Select Top-k locations for the optimization goals
29:         $rank_{tok}\leftarrow rep_{tok}[s_{tok}+n_{1}]$triangleright Get the replacement tokens on the selected locations
30:         $x^{\prime}_{p}\leftarrow\text{Swap}(rank_{tok},\text{input\_tokens},x_{p},x^{\prime}_{p},J_{2})$triangleright Alg. 3
@@
44:return $a,b,c,d$triangleright Final conditions
@@
1:A list of ranked tokens $rank$, input tokens, unperturbed question $x_{p}$, perturbed question $x^{\prime}_{p}$, and LLM Judge $J_{2}$ for checking the semantic correctness.
@@ A.1 Limitations
Our method requires significant computational resources due to the complexity of generating and evaluating successful examples on LLMs. The effectiveness of our approach relies on the LLM judge’s ability to assess the correctness of reasoning within CoT explanations. However, this evaluation process is not entirely reliable, as LLMs may exhibit inconsistencies or biases in their judgments, potentially affecting the accuracy of our robustness analysis. Further, the token-level attack may slightly impact the original meaning of the questions.
@@ A.2 Details of datasets
SingleEq [31] is a collection of algebraic word problems designed to evaluate the mathematical reasoning capabilities of machine learning models. Each problem in the dataset consists of a short natural language description corresponding to a single-variable linear equation, making it a benchmark for assessing symbolic reasoning and arithmetic problem-solving skills in language models.
GSM8K [9] is a large-scale benchmark designed to evaluate the mathematical reasoning abilities of language models. It comprises high-quality, grade-school-level arithmetic word problems, each requiring multi-step reasoning to derive the correct answer. The dataset was curated with a focus on problems that involve multi-step computations rather than simple numerical lookups or direct retrieval. GSM8K is widely used to assess Chain-of-Thought (CoT) prompting. It provides a challenging environment where intermediate reasoning steps are crucial for arriving at the correct solution. The dataset has become a standard benchmark in mathematical problem-solving and reasoning research, particularly in the context of LLM fine-tuning and evaluation.
StrategyQA [15] is a question-answering benchmark. Unlike traditional datasets, it focuses on questions where the necessary reasoning steps are implicit and must be inferred using a strategy. The dataset comprises 2780 examples, each including a strategy question, its decomposition into reasoning steps, and supporting evidence paragraphs. Analysis indicates that questions in StrategyQA are concise, cover diverse topics, and require a wide range of reasoning strategies. It is widely used in commonsense reasoning.
@@ A.4 Additional ablation studies
Embedding perturbation percentage $\epsilon$. In Figure 5, for embedding-level attacks, we show UR, WR, and SR on SingleEq, GSM8K, and StrategyQA datasets with varying perturbation percentage $\epsilon$ values in $[0.005,0.01,0.02,0.03,0.05]$ using Llama-3-8B. A larger epsilon leads to a smaller UR and a larger WR. Also, a tradeoff exists for finding the best $\epsilon$ for good SRs, which usually falls in the middle. According to each dataset’s result, we select $\epsilon=0.02,0.005,0.03$ for SingleEq, GSM8K, and StrategyQA respectively, with good SRs.

### v2

@@ 5.1 Experimental setup
Datasets. We use three datasets in line with existing literature on reasoning (Wei et al., 2022b; Trivedi et al., 2022; Kojima et al., 2022b; Diao et al., 2023). For math reasoning, we select two math datasets SingleEq (Koncel-Kedziorski et al., 2016) with single-step math problems and GSM8K (Cobbe et al., 2021) with multi-step math problems. Apart from that, we select StrategyQA (Geva et al., 2021) for commonsense reasoning. For all datasets, we use the test split for the evaluation. Specifically, we sampled $500$ questions for the evaluation of the StrategyQA dataset and used all data (508 for SingleEq and 1319 for GSM8K) in the other two datasets. More details about the datasets can be found in Appendix A.2.
Models. Our experiments focus on open-source and closed-source models for robustness evaluation. For open-source models, we use Llama-3-8B (Dubey et al., 2024a), Mistral-7B (Jiang et al., 2023), Zephyr-7B-beta (Tunstall et al., 2023), Qwen2.5-7B (Yang et al., 2024), and DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025b). We evaluate the transferability of white-box pertrubations to closed-source models GPT-4 and GPT-3.5-turbo (Achiam et al., 2023). Also, we use GPT-3.5-turbo as our judging model for comparing COTs, and DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025b) for evaluating the semantic meanings of perturbed questions. Human evaluations of those judges are available in Appendix A.4 and Appendix A.7.
@@
Implementation. For token-level perturbation on open-source models, we set the inserted ratio to be $a=0.2$ for all datasets; for the replacement ratio $k$, we set it to be $(0.5,0.25,0.5)$ for SingleEq, GSM8K, and StrategyQA datasets, respectively. For embedding-level perturbation, we choose hyperparameters with perturbation percentage $\epsilon=(0.02,0.005,0.03)$ for SingleEq, GSM8K, and StrategyQA datasets, which are imperceptible for the token mapping. Also, we set the number of perturbation steps $j$ to be 5 for all datasets and two kinds of perturbations. For closed-source model experiments, we set temperature $T=0.7$ and we test the closed-source models on the adversarial examples that open-source models generate. The default version of GPT-3.5-turbo used is GPT-3.5-turbo-0125, and the GPT-4o used is GPT-4o-2024-08-06. Apart from that, we set the maximum length of generating output to be 256. We use a conventional ICL setting for our experiments (Brown et al., 2020; Rae et al., 2021). We apply the same number of exemplars as Wei et al. (2022b) and use the exemplars from Diao et al. (2023), which are 8 for SingleEq and GSM8K, as well as 6 for StrategyQA. Runtime, LLM judge analysis, and limitations can be found in Appendix A.5 and A.7.
@@ 5.2 Comparison with baselines
In Table 3, we compare MATCHA with the following baselines for token-level perturbations: (1) Random: we randomly insert tokens into the original question; (2) GCG: we optimize a suffix similar to Zou et al. (2023) using our defined loss; (3,4,5) MATCHA: we experiment with MATCHA with random locations/with only CoT loss/full versions. We show that MATCHA can achieve a comparatively lower unattackable rate and higher success rate, indicating MATCHA token-level perturbations have stronger perturbation effectiveness.
@@ 5.4 Transferability to closed-source models
Model-level analysis: Qwen2.5-7B produces notably high transferability to GPT-4o, especially on structured tasks like SingleEq and GSM8K, indicating that its perturbations align closely with GPT-4o’s decision boundaries. Zephyr-7B and Mistral-7B show strong transferability to GPT-3.5-turbo on different tasks, suggesting that base model capabilities and reasoning biases shape adversarial generalization.
Dataset-level analysis: Tasks like StrategyQA exhibit high transferability to GPT-3.5-turbo but much lower transferability to GPT-4o, suggesting that commonsense reasoning attacks transfer less effectively to more advanced models like GPT-4o. In contrast, GSM8K shows more consistent transferability, likely due to its focus on multi-step math reasoning, which may yield more structurally generalizable perturbations.
These results underscore that both types of reasoning in the dataset and the characteristics of the base model jointly determine transferability.
@@ 5.5 Examples of successful attacks
Wrong calculations refers to the errors in the calculation steps. As shown in the first row of Figure 3, the model thinks "20/3=5", which is incorrect in the calculation. Wrong reasoning refers to the evident reasoning errors in the procedure of arriving at the correct solution. For the second example in Figure 3, the model makes the mistake of getting the total number of running hours (should be 3+2/32=63+2/3*2=6 instead of $4.5$), leading to the overall error in the reasoning procedures. Wrong information refers to the model outputting some information that is wrong according to the question, to arrive at the correct answer. We present an example in the third column of Figure 3, where the model makes up "Sophocles" instead of "Sophist" to arrive at the correct answer. Unrelated information refers to the model outputting some unrelated information when trying to answer the question. In the fourth column of Figure 3, the model starts to make up non-existent information like "There are 15 trees in the grove" to guess the answer. The complete list of successful attack examples can be found in Appendix A.6.
@@ 6 Conclusion
In this work, we empirically examine the Decoupling Hypothesis, questioning the assumption that correct answers in LLMs necessarily arise from faithful reasoning. Using our proposed MATCHA framework, we introduce Answer-Conditioned Probing to disentangle the robustness of a model’s prediction from that of its justification. We find that LLMs can often produce correct answers despite fragile or inconsistent reasoning. In particular, small, imperceptible perturbations can disrupt the logical structure of a Chain-of-Thought while leaving the final answer unchanged, especially in multi-step and commonsense tasks. The transferability of these effects to black-box models (e.g., GPT-4) suggests that this phenomenon is not model-specific, but reflects a broader limitation of current training and alignment practices. Overall, MATCHA provides a diagnostic perspective on reasoning robustness, highlighting the gap between answer correctness and process reliability. Our results suggest that improving trustworthiness will require moving beyond answer-level supervision toward methods that better couple reasoning structure with model predictions.
@@ Appendix A Appendix
1:Number of attack steps $k$, ICL examples II with length $n_{1}$, unlabeled question $x_{p}$ with length $n_{2}$, ground truth $y_{gt}$, model $f$ parameterized by $\theta$, attack region $\epsilon$, stepsize $\alpha$, loss weight $\lambda$, embedding layer $h$, judge LLM $J$.
@@
[…]
@@
10:  $E_{a}\leftarrow h(I,(x_{p},y_{p}))$ $\triangleright$ Get embeddings
11:  $q_{a},q_{c}\leftarrow f(E_{a}),f(E_{c})$ $\triangleright$ Get logits
@@
16:   $\mathcal{L}_{opt}\leftarrow\mathcal{L}_{c}-\lambda\cdot\mathcal{L}_{a}$ $\triangleright$ Loss objective
@@
19:   $E^{\prime}_{c}\leftarrow\text{concat}(E_{c}[:n_{1}],E_{tmp}[n_{1}:n_{2}],E_{c}[n_{2}:])$ $\triangleright$ Update perturbed embedding
@@
33:return $a,b,c,d$ $\triangleright$ Final conditions
@@
1:Number of attack steps $k$, ICL examples II with length $n_{1}$, input ids and input tokens, unlabeled question $x_{p}$ with length $n_{2}$, ground truth $y_{gt}$, model $f$ parameterized by $\theta$, embedding layer $h$, selected number of token $T_{k}$, judge LLM $J_{1}$ for reasoning correctness and $J_{2}$ for question correctness.
@@
[…]
@@
9:  $q_{a},q_{c}\leftarrow f(I,(x_{p},y_{p})),f(I,(x_{p},y_{p},r_{p}))$ $\triangleright$ Get target logits
@@
11:  $x^{\prime}_{r}\leftarrow\text{random\_insert}(x^{\prime}_{p})$ $\triangleright$ Randomly insert tokens into the question part
@@
20:  $\text{insert\_indices}\leftarrow\text{get\_insert\_indices}(x^{\prime}_{r},x^{\prime}_{p})$ $\triangleright$ Get inserted indices
@@
22:   $q^{\prime}_{c},q^{\prime}_{a}\leftarrow f(I,(x_{p},y_{p})),f(I,(x^{\prime}_{r},y_{p},r^{\prime}_{p}))$ $\triangleright$ Get perturbed logits
@@
24:   $\mathcal{L}_{opt}\leftarrow\mathcal{L}_{c}-\lambda\cdot\mathcal{L}_{a}$ $\triangleright$ Loss objective
25:   $g_{tok}\leftarrow\nabla_{e}(-\mathcal{L}_{opt})$ $\triangleright$ $g_{tok}$ has a shape of [seq_len, vocal_size]
26:   $rep_{tok}\leftarrow\text{Top1}(g_{tok},\text{axis}=1)[\text{insert\_indices}]$ $\triangleright$ We get the top 1 replacement token for each inserted token on the question part
27:   $s_{tok}\leftarrow\text{gather}(g_{tok},\text{input\_ids})[\text{insert\_indices}]$ $\triangleright$ Get gradients of inserted locations on the question part
28:   $s_{tok}\leftarrow\text{Topk}(s_{tok},T_{k})$ $\triangleright$ Select Top-k locations for the optimization goals
29:   $rank_{tok}\leftarrow rep_{tok}[s_{tok}+n_{1}]$ $\triangleright$ Get the replacement tokens on the selected locations
30:   $x^{\prime}_{p}\leftarrow\text{Swap}(rank_{tok},\text{input\_tokens},x_{p},x^{\prime}_{p},J_{2})$ $\triangleright$ Alg. 3
@@
44:return $a,b,c,d$ $\triangleright$ Final conditions
@@
1:A list of ranked tokens rankrank, input tokens, unperturbed question $x_{p}$, perturbed question $x^{\prime}_{p}$, and LLM Judge $J_{2}$ for checking the semantic correctness.
@@ A.1 Limitations
Our method requires significant computational resources due to the complexity of generating and evaluating successful examples on LLMs. The effectiveness of our approach relies on the LLM judge’s ability to assess the correctness of reasoning within CoT explanations. However, this evaluation process is not entirely reliable, as LLMs may exhibit inconsistencies or biases in their judgments, which could potentially affect the accuracy of our robustness analysis. Further, the token-level attack may slightly impact the original meaning of the questions.
@@ A.2 Details of datasets
SingleEq (Koncel-Kedziorski et al., 2016) is a collection of algebraic word problems designed to evaluate the mathematical reasoning capabilities of machine learning models. Each problem in the dataset consists of a short natural language description corresponding to a single-variable linear equation, making it a benchmark for assessing symbolic reasoning and arithmetic problem-solving skills in language models.
GSM8K (Cobbe et al., 2021) is a large-scale benchmark designed to evaluate the mathematical reasoning abilities of language models. It comprises high-quality, grade-school-level arithmetic word problems, each requiring multi-step reasoning to derive the correct answer. The dataset was curated with a focus on problems that involve multi-step computations rather than simple numerical lookups or direct retrieval. GSM8K is widely used to assess Chain-of-Thought (CoT) prompting. It provides a challenging environment where intermediate reasoning steps are crucial for arriving at the correct solution. The dataset has become a standard benchmark in mathematical problem-solving and reasoning research, particularly in the context of LLM fine-tuning and evaluation.
StrategyQA (Geva et al., 2021) is a question-answering benchmark. Unlike traditional datasets, it focuses on questions where the necessary reasoning steps are implicit and must be inferred using a strategy. The dataset comprises 2780 examples, each including a strategy question, its decomposition into reasoning steps, and supporting evidence paragraphs. Analysis indicates that questions in StrategyQA are concise, cover diverse topics, and require a wide range of reasoning strategies. It is widely used in commonsense reasoning.
@@ A.4 Additional ablation studies
Embedding perturbation percentage $\epsilon$. In Figure 5, for embedding-level attacks, we show UR, WR, and SR on SingleEq, GSM8K, and StrategyQA datasets with varying perturbation percentage $\epsilon$ values in $[0.005,0.01,0.02,0.03,0.05]$ using Llama-3-8B. A larger epsilon leads to a smaller UR and a larger WR. Also, a tradeoff exists for finding the best $\epsilon$ for good SRs, which usually falls in the middle. According to the results, we select $\epsilon=0.02,0.005,0.03$ for SingleEq, GSM8K, and StrategyQA respectively, with good SRs.
@@
Unnatural examples for token-level perturbations. We acknowledge that preserving the final answer alone does not guarantee that the semantic meaning of the question remains unchanged. To assess this directly, we performed an additional human evaluation to examine the semantic fidelity of the perturbed inputs. Concretely, we randomly sampled 50 examples from each dataset (SingleEq, GSM8K, and StrategyQA) and asked human annotators to judge whether the meaning of the perturbed question differed substantially from that of the original. Across all datasets, we observe a high degree of semantic preservation, with agreement rates of 98% on SingleEq, 94% on GSM8K, and 88% on StrategyQA (Table 4). These results indicate that, in the vast majority of cases, the perturbations do not materially alter the original question intent, supporting the validity of using LLM-based judges within our evaluation framework.
Dataset Agreement
SingleEq 98%
GSM8K 94%
StrategyQA 88%
Table 4: Human evaluation of semantic agreement between original and perturbed questions.
Full results with confidence intervals for Figure 4. The results are shown in Table 5 and Table 6. We observe that a ratio of 0.2 consistently performs well across datasets, with stable performance within the confidence bounds. Increasing the number of steps beyond 5 does not lead to notable improvements, and the differences fall within the margin of error. To ensure efficiency, we use 5 perturbation steps in our main experiments.
Token Insert Ratio Unattackable (Mean $\pm$ 95% CI) Success (Mean $\pm$ 95% CI)
0.1 0.696 $\pm$ 0.021 0.127 $\pm$ 0.021
0.2 0.451 $\pm$ 0.056 0.152 $\pm$ 0.056
0.3 0.422 $\pm$ 0.021 0.064 $\pm$ 0.021
0.4 0.456 $\pm$ 0.037 0.015 $\pm$ 0.037
Table 5: Effect of token insertion ratio on unattackable rate and attack success rate.
Steps Unattackable (Mean $\pm$ 95% CI) Success (Mean $\pm$ 95% CI)
5 0.338 $\pm$ 0.037 0.162 $\pm$ 0.037
10 0.338 $\pm$ 0.097 0.157 $\pm$ 0.076
15 0.441 $\pm$ 0.000 0.152 $\pm$ 0.042
Table 6: Effect of the number of perturbation steps on unattackable rate and attack success rate.